### v1

# An improvement toward global boundedness in a fully parabolic chemotaxis with singular sensitivity in any dimension

Abstract
We are concerned with the following chemotaxis system
$\begin{cases}u_{t}=\Delta u-\chi\nabla\cdot\left(\frac{u}{v}\nabla v\right),\\ v_{t}=\Delta v-u+v,\end{cases}$
in an open bounded domain with smooth boundary $\Omega\subset\mathbb{R}^{n}$ with $n\geq 3$. It is known that solutions are globally bounded when $0<\chi<\sqrt{\frac{2}{n}}$. In this paper, by considering the following energy functional
$F_{\lambda}(u,v)=\int_{\Omega}\frac{u^{p}}{v^{r}}+\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\int_{\Omega}v^{p-r}$
for $p>1$, $r=\frac{p-1}{2}$, and $\lambda>0$, we improve this result by showing that there exists $\chi_{0}>\sqrt{\frac{2}{n}}$ such that for any $\chi\in(0,\chi_{0})$, solutions exist globally and remain bounded.
Keywords: Chemotaxis, global boundedness, singular sensitivity
Mathematics Subject Classification: 35B35, 35K45, 35K55, 92C15, 92C17
1 Introduction
We consider the following system arising from a chemotaxis model in an open bounded domain $\Omega\subset\mathbb{R}^{n}$, where $n\geq 2$:
$\begin{cases}u_{t}=\Delta u-\chi\nabla\cdot\left(\frac{u}{v}\nabla v\right),\qquad&\text{in }\Omega\times(0,T_{\rm max}),\\ v_{t}=\Delta v-v+u,\qquad&\text{in }\Omega\times(0,T_{\rm max}),\end{cases}$ (1.1)
where $\chi>0$ and $T_{\rm max}\in(0,\infty]$ is the maximal existence time. The system is endowed with the homogeneous Neumann boundary condition
$\frac{\partial u}{\partial\nu}=\frac{\partial v}{\partial\nu}=0,\qquad(x,t)\in\partial\Omega\times(0,T_{\rm max}),$ (1.2)
and initial conditions
$\begin{cases}u_{0}\in C^{0}(\bar{\Omega})\text{ is nonnegative with }\int_{\Omega}u_{0}>0,\\ v_{0}\in W^{1,\infty}(\Omega),\quad v_{0}>0\quad\text{in }\bar{\Omega}.\end{cases}$ (1.3)
The system (1.1) was originally introduced in the seminal work [7], where it was proposed as a mathematical model to describe the phenomenon of chemotaxis—namely, the directed movement of cells or microorganisms toward or away from chemical signals in their environment. This mechanism plays a fundamental role in a variety of biological processes, such as the aggregation of cellular slime molds, immune response behavior, and the formation of patterns in bacterial populations. In the context of system (1.1), the function $u(x,t)$ typically denotes the density of a cell population (such as bacteria or amoebae), while $v(x,t)$ represents the concentration of a chemical substance that influences the movement of these cells at a given spatial location $x$ and time $t$. This interaction reflects the phenomenon of chemotaxis, where cellular motion is guided by gradients of chemical signals. A well-studied variant of this model arises when the second equation in system (1.1) is replaced by the elliptic equation $0=\Delta v-v+u$, leading to the so-called parabolic-elliptic system. Under the assumption of radially symmetric initial data, it was shown in [17] that solutions to this system exist globally in time whenever the chemotactic sensitivity parameter satisfies $\chi<\frac{2}{n-2}$, whereas solutions may blow up if $\chi>\frac{2n}{n-2}$. This result was later extended in [2], where the condition for global existence was relaxed to $\chi<\frac{2}{n}$ without requiring radial symmetry in the initial data. Moreover, global boundedness of solutions under this condition was established in [6]. In two spatial dimensions, a stronger result was obtained in [5], demonstrating that blow-up does not occur for any positive value of $\chi$, highlighting a critical difference between low and high-dimensional behavior. More recently, in [8], it was shown that there exists a threshold $\chi_{0}>\frac{2}{n}$ such that for all $\chi<\chi_{0}$, solutions are globally defined and remain uniformly bounded for all time. Despite these substantial contributions, the exact critical value of $\chi$ that delineates the transition between global existence and finite-time blow-up for the parabolic-elliptic system is still not fully characterized. For further advances related to parabolic-elliptic systems with singular sensitivity, we refer the reader to [23, 24, 13, 10, 9, 14].
We now turn our attention to the fully parabolic Keller–Segel system. It was established in [21] that classical solutions exist globally whenever $\chi<\sqrt{\frac{2}{n}}$, and later in [3] that such solutions remain uniformly bounded under the same condition. Besides, in [21], it was shown that global weak solutions exist when $\chi<\sqrt{\frac{n+2}{3n-4}}$. By defining a slightly more general concept of weak solutions, the authors in [18] proved the global existence of radial solutions when $\chi<\sqrt{\frac{n}{n-2}}$. Later, the construction of global generalized solutions was established in [12] under the restriction that
$\chi<\begin{cases}\infty\qquad\text{if }n=2,\\ \sqrt{8}\qquad\text{if }n=3,\\ \frac{n}{n-2}\qquad\text{if }n\geq 4.\end{cases}$
Therefore, these results naturally lead to the question of whether there exists a critical value $\chi_{c}>0$ such that solutions are globally bounded for all $\chi<\chi_{c}$, whereas for $\chi>\chi_{c}$ blow-up may occur. The result in [11] suggests that this critical threshold satisfies $\chi_{c}>1.016$ when $n=2$, indicating that $\chi=1$ is not the critical value in dimension two. Further, the authors in [4] conjectured that $\chi_{c}=\frac{n}{n-2}$, and provided a rigorous proof that solutions are global and bounded whenever $\chi<\frac{n}{n-2}$, assuming the left hand side of the first equation in system (1.1) is replaced by $\tau u_{t}$ with sufficiently small $\tau>0$. Without imposing smallness on $\tau$, the recent work [1] improved previous lower bound for $\chi_{c}$ in low dimensions by showing $\chi_{c}\geq\frac{4}{n}$. This result enhances those in [21] and [11] for dimensions $2\leq n\leq 8$, since $\frac{4}{n}\geq\sqrt{\frac{2}{n}}$ and $2>1.016$. In line with the result established in [4], it was shown in [22] that solutions are arbitrary large if $\tau$ is sufficiently small and $\chi>\frac{n}{n-2}$. This result clearly supports the aforementioned conjecture that $\chi_{c}=\frac{n}{n-2}$. Therefore, it is reasonable to conclude that the threshold $\chi_{c}$ exceeds $\sqrt{\frac{2}{n}}$, indicating that $\sqrt{\frac{2}{n}}$ is not the critical value in any dimension. This conclusion has been rigorously established for spatial dimensions up to eight or when $\tau$ is sufficiently small and the domain $\Omega$ is convex; however, whether it holds for dimensions $n\geq 9$, or without the smallness condition on $\tau$ and in nonconvex domains, remains an open problem. In this work, we address this gap by proving that $\chi_{c}>\sqrt{\frac{2}{n}}$ for all $n\geq 3$. Our main theorem is stated as follows:
Theorem 1.1.
Let $\Omega\subset\mathbb{R}^{n}$, with $n\geq 3$, be an open bounded domain with smooth boundary. Then there exists $\chi_{0}>\sqrt{\frac{2}{n}}$ such that, for any $\chi\in(0,\chi_{0})$, the system (1.1), with the boundary condition (1.2) and initial condition (1.3), admits a unique global positive classical solution $(u,v)\in\left[C^{2,1}\left(\bar{\Omega}\times(0,\infty)\right)\cap C^{0}\left(\bar{\Omega}\times[0,\infty)\right)\right]^{2}$ such that
$\sup_{t>0}\left\|u(\cdot,t)\right\|_{L^{\infty}(\Omega)}<\infty.$
Remark 1.1.
An explicit lower bound for $\chi_{0}$, established in the proof of Lemma 4.1, is $\sqrt{\frac{2(1+\delta)}{n}}$ with $\delta>\frac{(n-2)^{2}}{256(n-1)\left(16^{n}(3n^{2})^{n+1}+(912n^{2})^{n+1}\right)}$.
Remark 1.2.
The works [11], [4], and [1] all require the convexity of the domain $\Omega$, whereas Theorem 1.1 imposes no such condition.
In addition, we propose a novel approach for studying chemotaxis systems with singular sensitivity in arbitrary dimensions, drawing inspiration from the works [21], [11], and [19]. In [21], the global existence of solutions for $\chi<\sqrt{\frac{2}{n}}$ was established by analyzing the quantity $\int_{\Omega}u^{p}v^{-r}$ for $p>1$ and $r>0$. Subsequently, [11] suggested that the inclusion of an additional summand could yield improved results. In a different setting, [19] introduced an energy functional involving the term $\int_{\Omega}v^{-q+1}|\nabla v|^{q}$, which motivates us to consider the generalized form $\int_{\Omega}v^{-l}|\nabla v|^{2p}$. This leads to the identification of $\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}$ as a suitable additional term to the original expression $\int_{\Omega}u^{p}v^{-r}$. To further facilitate the analysis, we introduce the auxiliary term $\int_{\Omega}v^{p-r}$, which allows us to directly estimate the boundedness of $\int_{\Omega}u^{p}v^{-r}$ without relying on the iterative argument employed in [3, 16] and without requiring a prior bound on $\int_{\Omega}v^{p-r}$. Consequently, we construct the energy functional
$F_{\lambda}(u,v)=\int_{\Omega}u^{p}v^{-r}+\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\int_{\Omega}v^{p-r},\quad\text{with }r=\frac{p-1}{2},$
which enables us to establish an alternative proof of the global boundedness of solutions to (1.1). In particular, it can be shown—using Lemma 3.1—that $F_{0}(u,v)$ remains bounded whenever $\chi<\sqrt{\frac{2}{n}}$.
The structure of the paper is as follows: In Section 2, we recall the local well-posedness of and an extensibility property for solutions to (1.1) and establish several auxiliary inequalities that will be used in the subsequent analysis. Section 3 provides several identities and estimates for solutions. In the final section, we prove the time-independent $L^{p}$ boundedness of solutions and then complete the proof of the main result.
2 Preliminaries
Let us begin this section by recalling the local existence of and an extensibility criteria of solutions to the system (1.1) as established in [11][Theorem 2.3].
Lemma 2.1.
Let $n\geq 1$, $\Omega\subset\mathbb{R}^{n}$ a bounded, smooth domain and $q>n$. Then there exist $T_{\rm max}\in(0,\infty]$ and a unique pair of nonnegative functions $(u,v)$ such that
$u$ $\in C^{0}\left(\bar{\Omega}\times[0,T_{\rm max})\right)\cap C^{2,1}\left(\bar{\Omega}\times(0,T_{\rm max})\right)\quad\text{and }$
$v$ $\in C^{0}\left(\bar{\Omega}\times[0,T_{\rm max})\right)\cap C^{2,1}\left(\bar{\Omega}\times(0,T_{\rm max})\right)\cap L^{\infty}_{loc}\left([0,T_{\rm max});W^{1,q}(\Omega)\right),$
solving the system (1.1) with initial conditions (1.3) and boundary conditions (1.2). Moreover, $u>0$ and $v>0$ in $\bar{\Omega}\times(0,T_{\rm max})$, and
$\text{if }T_{\rm max}<\infty,\text{ then }\limsup_{t\to T_{\rm max}}\left\{\left\|u(\cdot,t)\right\|_{L^{\infty}(\Omega)}+\left\|v(\cdot,t)\right\|_{W^{1,q}(\Omega)}\right\}=\infty.$ (2.1)
From now on, we denote $(u,v)$ as a classical solution to (1.1) in $(0,T_{\rm max})$. The following lemma gives an uniform in time $L^{1}$ bound for $v$.
Lemma 2.2.
There exists positive constants $m$ and $C$ such that
$\int_{\Omega}u(\cdot,t)=m\qquad\text{for all }t\in(0,T_{\rm max}),$
and
$\int_{\Omega}v(\cdot,t)\leq C\qquad\text{for all }t\in(0,T_{\rm max}).$
Proof.
By integrating the first equation of (1.1) over $\Omega$, we obtain that
$\frac{d}{dt}\int_{\Omega}u(\cdot,t)=0\qquad\text{for all }t\in(0,T_{\rm max}),$
which entails that
$\int_{\Omega}u(\cdot,t)=m:=\int_{\Omega}u_{0}\qquad\text{for all }t\in(0,T_{\rm max}).$
Integrating the second equation of (1.1) over $\Omega$ yields
$\frac{d}{dt}\int_{\Omega}v+\int_{\Omega}v=m\qquad\text{for all }t\in(0,T_{\rm max}).$
Applying Gronwall’s inequality to this implies that
$\sup_{t\in(0,T_{\rm max})}\int_{\Omega}v(\cdot,t)\leq\max\left\{\int_{\Omega}v_{0},m\right\},$
which completes the proof. ∎
The next lemma gives us a time independent lower bound for $v$, which was established in [4][Lemma 3.1].
Lemma 2.3.
There exists a positive constant $\eta$ depending only on $\inf_{\bar{\Omega}}v_{0},\int_{\Omega}u_{0}$, and $\Omega$ such that
$\inf_{x\in\Omega}v(x,t)\geq\eta\qquad\text{for all }t\in(0,T_{\rm max}).$
The following lemma establishes regularity properties of $v$ based on the boundedness of $u$; for a detailed proof, we refer the reader to [21][Lemma 2.4].
Lemma 2.4.
Let $T\leq T_{\rm max}$ and $T<\infty$, and $1\leq\theta,\mu\leq\infty$.
  1. (i)
    If $\frac{n}{2}\left(\frac{1}{\theta}-\frac{1}{\mu}\right)<1$, then there exists $C>0$ independent of $T$ such that
    $\left\|v(\cdot,t)\right\|_{L^{\mu}(\Omega)}\leq C\left(1+\sup_{s\in(0,T)}\left\|u(\cdot,s)\right\|_{L^{\theta}(\Omega)}\right)\quad\text{for all }t\in(0,T).$
  2. (ii)
    If $\frac{1}{2}+\frac{n}{2}\left(\frac{1}{\theta}-\frac{1}{\mu}\right)<1$, then there exists $C>0$ independent of $T$ such that
    $\left\|\nabla v(\cdot,t)\right\|_{L^{\mu}(\Omega)}\leq C\left(1+\sup_{s\in(0,T)}\left\|u(\cdot,s)\right\|_{L^{\theta}(\Omega)}\right)\quad\text{for all }t\in(0,T).$
Now we will derive several useful inequalities similar to [20][Lemma 3.3] and [19][Lemma 3.4], which will be later applied in Lemma 3.3.
Lemma 2.5.
Let $l>0$ and $q>l+1$. For any $\phi\in C^{2}(\bar{\Omega})$ such that $\phi>0$ in $\bar{\Omega}$ and $\frac{\partial\phi}{\partial\nu}=0$ on $\partial\Omega$, the following inequalities hold
$\int_{\Omega}\phi^{-l-2}|\nabla\phi|^{q+2}\leq\left(\frac{q+\sqrt{n}}{l+1}\right)^{2}\int_{\Omega}\phi^{-l}|\nabla\phi|^{q-2}|D^{2}\phi|^{2},$ (2.2)
and
$\int_{\Omega}\phi^{-l-2}|\nabla\phi|^{q+2}\leq\left(4\left(\frac{l+1+\sqrt{n}}{q-l+1}\right)^{2}+2\right)\int_{\Omega}\phi^{-l+2}|\nabla\phi|^{q-2}|D^{2}\ln\phi|^{2},$ (2.3)
as well as
$\int_{\Omega}\phi^{-l}|\nabla\phi|^{q-2}|D^{2}\phi|^{2}\leq\left(\frac{q-l+1}{q-l-1}\cdot\left(\frac{l+1+\sqrt{n}}{q-l+1}\right)^{2}+1\right)\int_{\Omega}\phi^{-l+2}|\nabla\phi|^{q-2}|D^{2}\ln\phi|^{2}.$ (2.4)
Proof.
Setting
$A$ $=\int_{\Omega}\phi^{-l-2}|\nabla\phi|^{q+2},$
$B$ $=\int_{\Omega}\phi^{-l}|\nabla\phi|^{q-2}|D^{2}\phi|^{2},$
$C$ $=\int_{\Omega}\phi^{-l+2}|\nabla\phi|^{q-2}|D^{2}\ln\phi|^{2},$
$D$ $=\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q-2}\nabla\phi\cdot D^{2}\phi\cdot\nabla\phi.$
and employing integration by parts yields that
$A$ $=-\int_{\Omega}\nabla\cdot\left(\phi^{-l-2}|\nabla\phi|^{q}\nabla\phi\right)\phi$
$=(l+2)\int_{\Omega}\phi^{-l-2}|\nabla\phi|^{q+2}-\int_{\Omega}\phi^{-l-1}\nabla|\nabla\phi|^{q}\cdot\nabla\phi-\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q}\Delta\phi.$
By using the identity $\nabla|\nabla\phi|^{2}=2D^{2}\phi\cdot\nabla\phi$ and the inequality $|\Delta\phi|\leq\sqrt{n}|D^{2}\phi|$, we obtain that
$(l+1)A$ $=q\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q-2}\nabla\phi\cdot D^{2}\phi\cdot\nabla\phi+\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q}\Delta\phi$
$\leq(q+\sqrt{n})\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q}|D^{2}\phi|$
$\leq(q+\sqrt{n})\sqrt{AB},$
which further implies (2.2). We use integration by parts and Hölder’s inequality to derive (2.3) by estimating
$A$ $=\int_{\Omega}|\nabla\ln\phi|^{l+1}|\nabla\phi|^{q-l-1}\nabla\ln\phi\cdot\nabla\phi$
$=-\int_{\Omega}\nabla|\nabla\ln\phi|^{l+1}\cdot\nabla\ln\phi|\nabla\phi|^{q-l-1}\phi-\int_{\Omega}\phi|\nabla\ln\phi|^{l+1}|\nabla\phi|^{q-l-1}\Delta\ln\phi$
$\quad-\int_{\Omega}\phi|\nabla\ln\phi|^{l+1}\nabla\ln\phi\cdot\nabla|\nabla\phi|^{q-l-1}$
$=-(l+1)\int_{\Omega}\phi^{-l}|\nabla\phi|^{q-2}\nabla\phi\cdot D^{2}\ln\phi\cdot\nabla\phi-\int_{\Omega}\phi^{-l}|\nabla\phi|^{q}\Delta\ln\phi$
$\quad-(q-l-1)\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q-2}\nabla\phi\cdot D^{2}\phi\cdot\nabla\phi$
$\leq(l+1+\sqrt{n})\int_{\Omega}\phi^{-l}|\nabla\phi|^{q}|D^{2}\ln\phi|-(q-l-1)D$
$\leq(l+1+\sqrt{n})\sqrt{AC}-(q-l-1)D.$ (2.5)
Employing the identity
$|D^{2}\phi|^{2}=\phi^{2}|D^{2}\ln\phi|^{2}+\frac{2}{\phi}\nabla\phi\cdot\left(D^{2}\phi\cdot\nabla\phi\right)-\frac{|\nabla\phi|^{4}}{\phi^{2}},$
implies that
$B=C+2D-A.$ (2.6)
Combining (2) and (2.6) with straightforward simplifications deduces that
$A+\frac{q-l-1}{q-l+1}B\leq\frac{2(l+1+\sqrt{n})}{q-l+1}\sqrt{AC}+\frac{q-l-1}{q-l+1}C.$ (2.7)
Dropping the second term of the right hand side and using Young’s inequality entails that
$A\leq\frac{A}{2}+2\left(\frac{l+1+\sqrt{n}}{q-l+1}\right)^{2}C+C.$
This leads to
$A\leq\left(4\left(\frac{l+1+\sqrt{n}}{q-l+1}\right)^{2}+2\right)C,$ (2.8)
which proves (2.3). To prove (2.4), we apply Young’s inequality to (2.7) to obtain that
$A+\frac{q-l-1}{q-l+1}B\leq A+\left(\frac{l+1+\sqrt{n}}{q-l+1}\right)^{2}C+\frac{q-l-1}{q-l+1}C.$
This further leads to
$B\leq\left(\frac{q-l+1}{q-l-1}\cdot\left(\frac{l+1+\sqrt{n}}{q-l+1}\right)^{2}+1\right)C,$
which deduces (2.4). The proof is now complete.
The following lemma provides a bound for the boundary integral that arises in the analysis of the quantity $\frac{d}{dt}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}$, as encountered in Lemma 3.2. This allows us to work on a general domain $\Omega$ without the convexity assumption.
Lemma 2.6.
Let $p\geq 1$, $0<r<p$, and $\eta>0$. Then there exists $C=C(\eta,p,r,n,\Omega)>0$ such that the following holds
$\int_{\partial\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}\cdot\frac{\partial|\nabla\phi|^{2}}{\partial\nu}\leq\eta\int_{\Omega}\phi^{-p-r+2}|\nabla\phi|^{2p-2}|D^{2}\ln\phi|^{2}+C\int_{\Omega}\phi^{p-r},$ (2.9)
for any $\phi\in C^{2}(\bar{\Omega})$ which is such that $\phi>0$ in $\bar{\Omega}$ and $\frac{\partial\phi}{\partial\nu}=0$ on $\partial\Omega$.
Proof.
Thanks to [15][Lemma 4.2] and Trace Sobolev embedding from $W^{1,1}(\Omega)$ into $L^{1}(\partial\Omega)$, one can find $c_{1}=c_{1}(\Omega)>0$ and $c_{2}>0$ such that
$\frac{\partial|\nabla\phi|^{2}}{\partial\nu}\leq c_{1}|\nabla\phi|^{2}\quad\text{on }\partial\Omega,$
and
$\int_{\partial\Omega}|w|\leq c_{2}\int_{\Omega}|\nabla w|+c_{2}\int_{\Omega}|w|\quad\text{for all }w\in C^{1}(\bar{\Omega}).$
Therefore, we obtain that
$\int_{\partial\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}\cdot\frac{\partial|\nabla\phi|^{2}}{\partial\nu}$ $\leq c_{1}\int_{\partial\Omega}\phi^{-p-r}|\nabla\phi|^{2p}$
$\leq c_{1}c_{2}\int_{\Omega}\left|\phi^{-p-r}|\nabla\phi|^{2p}\right|+c_{1}c_{2}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p}.$ (2.10)
Using the identity $\nabla|\nabla\phi|^{2}=2D^{2}\phi\cdot\nabla\phi$ and Young’s inequality implies that
$c_{1}c_{2}\int_{\Omega}\left|\phi^{-p-r}|\nabla\phi|^{2p}\right|$ $\leq 2pc_{1}c_{2}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-1}|D^{2}\phi|+c_{1}c_{2}(p+r)\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}$
$\leq\epsilon\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}|D^{2}\phi|^{2}+\frac{p^{2}c_{1}^{2}c_{2}^{2}}{\epsilon}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p}$
$\quad+c_{1}c_{2}(p+r)\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}.$ (2.11)
Abbreviating $c_{3}:=c_{1}c_{2}+\frac{p^{2}c_{1}^{2}c_{2}^{2}}{\epsilon}$ and $c_{4}:=c_{1}c_{2}(p+r)$, and applying Young’s inequality with $\epsilon>0$ deduces that
$c_{3}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p}$ $=\int_{\Omega}\left\{\frac{\epsilon}{2}\phi^{-p-r-2}|\nabla\phi|^{2p+2}\right\}^{\frac{p}{p+1}}\cdot\left\{\left(\frac{2}{\epsilon}\right)^{\frac{p}{p+1}}c_{3}\phi^{\frac{p-r}{p+1}}\right\}$
$\leq\frac{\epsilon}{2}\int_{\Omega}\phi^{-p-r-2}|\nabla\phi|^{2p+2}+\frac{2^{p}}{\epsilon^{p}}c_{3}^{p+1}\int_{\Omega}\phi^{p-r},$ (2.12)
and
$c_{4}\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}$ $=\int_{\Omega}\left\{\frac{\epsilon}{2}\phi^{-p-r-2}|\nabla\phi|^{2p+2}\right\}^{\frac{2p+1}{2p+2}}\cdot\left\{\left(\frac{2}{\epsilon}\right)^{\frac{2p+1}{2p+2}}c_{4}\phi^{\frac{p-r}{2p+2}}\right\}$
$\leq\frac{\epsilon}{2}\int_{\Omega}\phi^{-p-r-2}|\nabla\phi|^{2p+2}+\left(\frac{2}{\epsilon}\right)^{2p+1}c_{4}^{2p+2}\int_{\Omega}\phi^{p-r}.$ (2.13)
Combining estimates from (2) to (2) entails that
$\int_{\partial\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}\cdot\frac{\partial|\nabla\phi|^{2}}{\partial\nu}\leq\epsilon\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}|D^{2}\phi|^{2}+\epsilon\int_{\Omega}\phi^{-p-r-2}|\nabla\phi|^{2p+2}+c_{5}\int_{\Omega}\phi^{p-r},$ (2.14)
where $c_{5}=\frac{2^{p}}{\epsilon^{p}}c_{3}+\left(\frac{2}{\epsilon}\right)^{2p+1}c_{4}^{2p+2}$. Thanks to Lemma 2.5, one can find $c_{6}=c_{6}(p,r,n)>0$ such that
$\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}|D^{2}\phi|^{2}+\int_{\Omega}\phi^{-p-r-2}|\nabla\phi|^{2p+2}\leq c_{6}\int_{\Omega}\phi^{-p-r+2}|\nabla\phi|^{2p-2}|D^{2}\ln\phi|^{2}$
Choosing $\epsilon=\frac{\eta}{c_{6}}$ and substituting this into (2.14) proves (2.9), thereby completing the proof.
3 A priori estimates
In this section, we establish several estimates for solutions to (1.1), which will be later applied to derive the most essential estimate, Lemma 4.1, allowing to obtain an $L^{\frac{n}{2}+}$ bound for $u$. Let us commence with the following lemma enabling us to control the quantity $\int_{\Omega}v^{p-r}$.
Lemma 3.1.
For any $p>1$, $10<r<p-$ 1, $M>0$ and $\epsilon>0$, there exists $C=C(p,r,M,\epsilon,u_{0},v_{0})>0$ such that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+M\int_{\Omega}v^{p-r}\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+C,$
for all $t\in(0,T_{\rm max})$.
Proof.
By testing the second equation of (1.1) by $(p-r)v^{p-r-1}$, we obtain that
$\frac{d}{dt}\int_{\Omega}v^{p-r}$ $=-(p-r)(p-r-1)\int_{\Omega}v^{p-r-2}|\nabla v|^{2}-(p-r)\int_{\Omega}v^{p-r}+(p-r)\int_{\Omega}uv^{p-r-1}.$ (3.1)
Applying Young’s inequality with $\epsilon>0$ deduces that
$(p-r)\int_{\Omega}uv^{p-r-1}$ $=\int_{\Omega}\left\{\epsilon\frac{u^{p+1}}{v^{r+1}}\right\}^{\frac{1}{p+1}}\cdot\left\{\epsilon^{-\frac{1}{p}}(p-r)^{\frac{p+1}{p}}v^{p-r}\right\}^{\frac{p}{p+1}}$
$\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{p-r}}+c_{1}\int_{\Omega}v^{p-r},$
where $c_{1}=\epsilon^{-\frac{1}{p}}(p-r)^{\frac{p+1}{p}}$. Thanks to Gagliardo–Nirenberg interpolation inequality, Young’s inequality, and Lemma 2.2 we obtain that
$\left(M+c_{1}-\frac{p+1}{2}\right)_{+}\int_{\Omega}v^{p-r}$ $\leq c_{2}\left(\int_{\Omega}|\nabla v^{\frac{p-r}{2}}|^{2}\right)^{\theta}\left(\int_{\Omega}v\right)^{(p-r)(1-\theta)}+c_{2}\left(\int_{\Omega}v\right)^{p-r}$
$\leq(p-r)(p-r-1)\int_{\Omega}v^{p-r-2}|\nabla v|^{2}+c_{3},$ (3.2)
where $\theta=\frac{\frac{p-r}{2}-\frac{1}{2}}{\frac{p-r}{2}+\frac{1}{n}-\frac{1}{2}}\in(0,1)$, $c_{2}>0$, and $c_{3}>0$. Finally, combining (3.1) and (3) proves the desired result. ∎
Lemma 3.2.
Let $p>1$ and $r=\frac{p-1}{2}$. Then there exists $C=C(p,n)>0$ such that the following holds
$\frac{d}{dt}\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$ $\leq-p\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}-(p-r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad-\frac{p-1}{4}\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+C\int_{\Omega}v^{p-r}$
$\quad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\quad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$ (3.3)
for any $t\in(0,T_{\rm max})$.
Proof.
Direct calculations show that
$\frac{d}{dt}\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$ $=2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}\nabla v\cdot\nabla v_{t}-(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}v_{t}$
$=I+J.$ (3.4)
By using of the second equation of (1.1) as well as the identities $\nabla v\cdot\nabla\Delta v=\frac{1}{2}\Delta|\nabla v|^{2}-|D^{2}v|^{2}$ and $\nabla|\nabla v|^{2}=2D^{2}v\cdot\nabla v$, and several integration by parts, we obtain that
$I$ $=2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}\nabla v\cdot\nabla\left(\Delta v-v+u\right)$
$=2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}\nabla v\cdot\nabla\Delta v-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p}+2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}\nabla v\cdot\nabla u$
$=p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}\Delta|\nabla v|^{2}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}|D^{2}v|^{2}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad+2p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u-2p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla|\nabla v|^{2}\cdot\nabla v$
$\quad-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v$
$=-p(p-1)\int_{\Omega}v^{-p-r}|\nabla v|^{2p-4}|\nabla|\nabla v|^{2}|^{2}+p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p-2}\nabla v\cdot\nabla|\nabla v|^{2}$
$\quad+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}|D^{2}v|^{2}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad+2p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u-4p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla v\cdot D^{2}v\cdot\nabla v$
$\quad-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v,$ (3.5)
and
$J$ $=-(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}\Delta v+(p+r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}-(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$
$=-(p+r)(p+r+1)\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p-2}\nabla|\nabla v|^{2}\cdot\nabla v$
$\quad+(p+r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}-(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u.$ (3.6)
Thanks to (3), (3), (3) and the identity that
$|D^{2}v|^{2}=v^{2}|D^{2}\ln v|^{2}+\frac{1}{v}\nabla v\cdot\nabla|\nabla v|^{2}-\frac{|\nabla v|^{4}}{v^{2}},$
we can write
$\frac{d}{dt}\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$ $=-2p\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}-p(p-1)\int_{\Omega}v^{-p-r}|\nabla v|^{2p-4}|\nabla|\nabla v|^{2}|^{2}$
$\quad+2p(p+r-1)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p-2}\nabla|\nabla v|^{2}\cdot\nabla v-(p-r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad-\left[(p+r)(p+r+1)-2p\right]\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$
$\quad-4p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla v\cdot D^{2}v\cdot\nabla v-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v$
$\quad+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}.$
Applying the inequality that $|\Delta v|\leq\sqrt{n}|D^{2}v|$ and Young’s inequality yields that
$\frac{d}{dt}\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$ $\leq-2p\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}-(p-r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad-\left[(p+r)(p+r+1)-2p-\frac{p(p+r-1)^{2}}{p-1}\right]\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}$
$\quad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\quad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}.$ (3.7)
Using Lemma 2.6 and noting that $(p+r)(p+r+1)-2p-\frac{p(p+r-1)^{2}}{p-1}=\frac{p-1}{4}$ since $r=\frac{p-1}{2}$, (3) deduces
$\frac{d}{dt}\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$ $\leq-p\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}-(p-r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad-\frac{p-1}{4}\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+c_{1}\int_{\Omega}v^{p-r}$
$\quad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\quad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u,$ (3.8)
for some $c_{1}>0$. The proof is now complete.
The following lemma, which justifies the consideration of the energy functional $F_{\lambda}(u,v)$, indicates that the quantity $\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}$ can be controlled by $\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}$.
Lemma 3.3.
Let $p\in\left(\frac{n}{2},n\right)$ and $r=\frac{p-1}{2}$. Then there exist positive constants $C=C(p,n)>0$ and $L_{1}$, $L_{2}$ and $L_{3}$ depending only on $n$ such that
$\frac{d}{dt}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+L_{1}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+L_{2}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}\leq L_{3}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+C\int_{\Omega}v^{p-r}$ (3.9)
for all $t\in(0,T_{\rm max})$.
Proof.
Using Lemma 3.2 and noting that $\frac{n}{2}<p<n$ yields that
$\frac{d}{dt}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}$ $+\frac{n+2}{8}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\frac{n-2}{8}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+\frac{n}{2}\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}$
$\leq c_{1}\int_{\Omega}v^{p-r}+c_{2}\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|+c_{3}\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u,$ (3.10)
where $c_{1}=c_{1}(p,n)>0$, $c_{2}=2n(2n-2+\sqrt{n})$, and $c_{3}=\frac{(2n-1)(3n-1)}{2}$. Thanks to Lemma 2.5, we obtain that
$\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}$ $\leq\left(4\left(\frac{p+r+1+\sqrt{n}}{p-r+1}\right)^{2}+2\right)\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}$
$\leq 102\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2},$ (3.11)
where the last inequality holds due to
$\frac{p+r+1+\sqrt{n}}{p-r+1}=\frac{3p+1+2\sqrt{n}}{p+3}<5.$ (3.12)
Similarly, we also have
$\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}|D^{2}v|^{2}$ $\leq\left(\frac{p-r+1}{p-r-1}\left(\frac{p+r+1+\sqrt{n}}{p-r+1}\right)^{2}+1\right)\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}$
$\leq 226\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2},$ (3.13)
where the last inequality is true because of (3.12) and
$\frac{p-r+1}{p-r-1}=\frac{p+3}{p-1}<9.$
In light of Young’s inequality, we obtain that
$c_{3}\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$ $\leq\frac{n-2}{16}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+\left(\frac{16}{n-2}\right)^{p}c_{3}^{p+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}$
$\leq\frac{n-2}{16}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+16^{n}c_{3}^{n+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}.$ (3.14)
Applying Hölder’s inequality, (3), (3) and Young’s inequality entails that
$c_{2}\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$ $\leq c_{2}\left(\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}|D^{2}v|^{2}\right)^{\frac{1}{2}}\left(\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}\right)^{\frac{p-1}{2(p+1)}}\left(\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}\right)^{\frac{1}{p+1}}$
$\leq 152c_{2}\left(\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}\right)^{\frac{p}{p+1}}\left(\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}\right)^{\frac{1}{p+1}}$
$\leq\frac{n}{2}\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}+(152c_{2})^{n+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}.$ (3.15)
Finally, combining (3), (3) and (3) deduces (3.9) with $L_{1}=\frac{n+2}{8}$, $L_{2}=\frac{n-2}{16}$ and $L_{3}=16^{n}c_{3}^{n+1}+(152c_{2})^{n+1}$. ∎
The following identity plays a crucial role in establishing uniform boundedness in $L^{p}(\Omega)$; for a detailed proof, we refer the reader to [21][Lemma 2.3].
Lemma 3.4.
Let $p\in\mathbb{R}$ and $r\in\mathbb{R}$. Then the following identity hold for any $t\in(0,T_{\rm max})$:
$\frac{d}{dt}\int_{\Omega}u^{p}v^{-r}+r\int_{\Omega}u^{p+1}v^{-r-1}$ $=-p(p-1)\int_{\Omega}u^{p-2}v^{-r}|\nabla u|^{2}-\left[r(r+1)+pr\chi\right]\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}$
$\quad+\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v+r\int_{\Omega}u^{p}v^{-r}.$
As a consequence, we have the following key estimate which will play an important role in deriving the boundedness of $F_{\lambda}(u,v)$.
Lemma 3.5.
There exist positive constants $L_{4}$ and $L_{5}$ depending only on $n$ such that for any $p\in\left(\frac{n}{2},n\right)$ the following holds
$\frac{d}{dt}\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}u^{p}v^{-r}+L_{4}\int_{\Omega}u^{p+1}v^{-r-1}\leq L_{5}(\chi^{2}p-1)_{+}\int_{\Omega}u^{p}v^{-p-r-2}|\nabla v|^{2}+C\int_{\Omega}v^{p-r},$ (3.16)
where $r=\frac{p-1}{2}$ and $C=C(p,n)>0$.
Proof.
From Lemma 3.4, we have that
$I:=\frac{d}{dt}\int_{\Omega}u^{p}v^{-r}+r\int_{\Omega}u^{p+1}v^{-r-1}$ $=-p(p-1)\int_{\Omega}u^{p-2}v^{-r}|\nabla u|^{2}-\left[r(r+1)+pr\chi\right]\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}$
$\quad+\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v+r\int_{\Omega}u^{p}v^{-r}.$ (3.17)
We apply Young’s inequality to obtain that
$\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v$ $\leq p(p-1)\int_{\Omega}u^{p-2}v^{-r}|\nabla u|^{2}$
$\quad+\frac{\left[2pr+p(p-1)\chi\right]^{2}}{4p(p-1)}\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}.$
Therefore, (3) yields
$I$ $\leq-h(p.r,\chi)\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}+r\int_{\Omega}u^{p}v^{-r},$ (3.18)
where
$h(p,r,\chi)=r(r+1)+pr\chi-\frac{\left[2pr+p(p-1)\chi\right]^{2}}{4p(p-1)}=-\frac{(p-1)(\chi^{2}p-1)}{4}.$
By Young’s inequality, one can find $c_{1}=c_{1}(p)>0$ such that
$(r+1)\int_{\Omega}u^{p}v^{-r}\leq\frac{r}{2}\int_{\Omega}u^{p+1}v^{-r-1}+c_{1}\int_{\Omega}v^{p-r}.$ (3.19)
Collecting (3), (3.18) and (3.19) and noting that $\frac{r}{2}>\frac{n-2}{8}$ deduces (3.16) with $L_{4}=\frac{n-2}{8}$ and $L_{5}=\frac{n-1}{4}$, which completes the proof.
4 Proof of the main result
Let us begin this section by showing that the energy function $F_{\lambda}(u,v)$ with an appropriate choice for $\lambda$ is uniformly bounded.
Lemma 4.1.
For any $n\geq 3$, there exists $\delta=\delta(n)>0$ such that for any $0<\chi<\sqrt{\frac{2(1+\delta)}{n}}$ and $1<p<\min\left\{\frac{1+\delta}{\chi^{2}},n\right\}$, there exists a constant $C=C(u_{0},v_{0},n,\Omega,p)>0$ such that
$\int_{\Omega}v^{p-r}+\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}v^{-p-r}|\nabla v|^{2p}\leq C,\qquad\text{for all }t\in(0,T_{\rm max}),$ (4.1)
where $r=\frac{p-1}{2}$.
Proof.
We assume first that $p>\frac{n}{2}$, and define $y(t):=\int_{\Omega}u^{p}v^{-r}+\lambda\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$, where $\lambda=\frac{L_{4}}{2L_{3}}$. From Lemma 3.3 and Lemma 3.5, we find that
$y^{\prime}(t)+c_{1}y(t)+\frac{L_{4}L_{2}}{2L_{3}}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+\frac{L_{4}}{2}\int_{\Omega}u^{p+1}v^{-r-1}$ $\leq L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}$
$\quad+c_{2}\int_{\Omega}v^{p-r}.$ (4.2)
$c_{1}=\min\left\{1,\frac{L_{1}L_{4}}{2L_{3}}\right\}$ and $c_{2}>0$. Applying Young’s inequality entails that
$L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}$ $=L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}\frac{u^{p}}{v^{\frac{p(r+1)}{p+1}}}\cdot\frac{|\nabla v|^{2}}{v^{\frac{p+r+2}{p+1}}}$
$\leq L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}u^{p+1}v^{-r-1}+L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}.$ (4.3)
Denoting $\delta=\min\left\{\frac{L_{2}L_{4}}{2L_{3}L_{5}},\frac{L_{4}}{2L_{5}}\right\}>\frac{(n-2)^{2}}{256(n-1)\left(16^{n}(3n^{2})^{n+1}+(912n^{2})^{n+1}\right)}$, then for any $\frac{n}{2}<p<\frac{1+\delta}{\chi^{2}}$ we have that $L_{5}\left(\chi^{2}p-1\right)_{+}\leq\min\left\{\frac{L_{2}L_{4}}{2L_{3}},\frac{L_{4}}{4}\right\}$. This, together with (4) and (4) implies that
$y^{\prime}(t)+c_{1}y(t)+\frac{L_{4}}{4}\int_{\Omega}u^{p+1}v^{-r-1}\leq c_{2}\int_{\Omega}v^{p-r},\qquad\text{for all }t\in(0,T_{\rm max}).$ (4.4)
Applying Lemma 3.1 with $M=c_{2}+c_{1}$ and $\epsilon=\frac{L_{4}}{4}$ entails that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+(c_{2}+c_{1})\int_{\Omega}v^{p-r}\leq\frac{L_{4}}{4}\int_{\Omega}u^{p+1}v^{-r-1}+c_{3},$ (4.5)
for some $c_{3}>0$. Combining (4.4) and (4.5) implies that
$\frac{d}{dt}F_{\lambda}(u,v)+c_{1}F_{\lambda}(u,v)\leq c_{3}\qquad\text{for all }t\in(0,T_{\rm max}),$
which combines with Gronwall’s inequality deduces (4.1) for any $\frac{n}{2}<p<\frac{1+\delta}{\chi^{2}}$. Now, for any $1<q<p$, applying Hölder’s inequality and Lemma 2.2 entails that
$\int_{\Omega}u^{q}v^{-\frac{q-1}{2}}+\int_{\Omega}v^{-\frac{3q-1}{2}}|\nabla v|^{2q}$ $\leq\left(\int_{\Omega}u^{p}v^{-\frac{p-1}{2}}\right)^{\frac{q}{p}}\left(\int_{\Omega}v^{\frac{1}{2}}\right)^{1-\frac{q}{p}}+\left(\int_{\Omega}v^{-\frac{3p-1}{2}}|\nabla v|^{2p}\right)^{\frac{q}{p}}\left(\int_{\Omega}v^{\frac{1}{2}}\right)^{1-\frac{q}{p}}$
$\leq|\Omega|^{\frac{1}{2}-\frac{q}{2p}}\left(\int_{\Omega}v\right)^{\frac{1}{2}-\frac{q}{2p}}\left\{\left(\int_{\Omega}u^{p}v^{-\frac{p-1}{2}}\right)^{\frac{q}{p}}+\left(\int_{\Omega}v^{-\frac{3p-1}{2}}|\nabla v|^{2p}\right)^{\frac{q}{p}}\right\}.$
Therefore, (4.1) also holds for any $1<p\leq\frac{n}{2}$. The proof is now complete. ∎
As a consequence, we can now derive an $L^{\frac{n}{2}+}$ for $u$ as follows:
Corollary 4.2.
Assume that $\chi<\sqrt{\frac{2(1+\delta)}{n}}$ where $\delta$ is given in Lemma 4.1. Then there exists $p>\frac{n}{2}$ such that
$\sup_{t\in(0,T_{\rm max})}\int_{\Omega}u^{p}(\cdot,t)<\infty.$
Proof.
Applying Lemma 4.1 yields that there exists $p_{0}\in\left(\frac{n}{2},n\right)$ such that
$\sup_{t\in(0,T_{\rm max})}\int_{\Omega}u^{p_{0}}v^{-r}:=c_{1}<\infty,$ (4.6)
where $r=\frac{p_{0}-1}{2}$. For any $T<T_{\rm max}$, setting
$M(T):=\sup_{t\in(0,T)}\int_{\Omega}u^{p}(\cdot,t)$
then we see from Lemma 2.1 that $M(T)<\infty$. Since $p_{0}>\frac{n}{2}$ implies that $\frac{p_{0}+1}{2(n-p_{0}+1)}>\frac{n}{2}$, we can choose $p\in\left(\frac{n}{2},\min\left\{p_{0},\frac{p_{0}+1}{2(n-p_{0}+1)}\right\}\right)$. Employing Hölder’s inequality entails that
$\int_{\Omega}u^{p}\leq\left(\int_{\Omega}u^{p_{0}}v^{-r}\right)^{\frac{p}{p_{0}}}\cdot\left(\int_{\Omega}v^{\frac{pr}{p_{0}-p}}\right)^{\frac{p_{0}-p}{p_{0}}},$
which combines with (4.6) implies that
$\left\|u(\cdot,t)\right\|_{L^{p}(\Omega)}\leq c_{1}^{\frac{1}{p_{0}}}\left\|v(\cdot,t)\right\|_{L^{\frac{pr}{p_{0}-p}}(\Omega)}^{\frac{r}{p_{0}}}\qquad\text{for all }t\in(0,T_{\rm max}).$ (4.7)
Since $\frac{n}{2}\left(\frac{1}{p}-\frac{p_{0}-p}{pr}\right)<1$ when $p<\frac{p_{0}+1}{2(n-p_{0}+1)}$, we apply Lemma 2.4 to obtain that
$\sup_{t\in(0,T)}\left\|v(\cdot,t)\right\|_{L^{\frac{pr}{p_{0}-p}}(\Omega)}\leq c_{2}\left(1+\sup_{t\in(0,T)}\int_{\Omega}u^{p}(\cdot,t)\right),$
for some $c_{2}>0$ independent of $T$. This, conjunction with (4.7) deduces that
$M(T)\leq c_{3}\left(1+M(T)^{\frac{r}{p_{0}}}\right)\qquad\text{for any }T<T_{\rm max},$
for some $c_{3}>0$ independent of $T$. Since $\frac{r}{p_{0}}<1$, this entails that $M(T)\leq c_{4}$, for some $c_{4}>0$ independent of $T$, which further deduces that $u\in L^{\infty}\left((0,T_{\rm max});L^{p}(\Omega)\right)$. ∎
The following lemma enables us to derive $L^{\infty}(\Omega\times(0,T_{\rm max}))$ bounds for the solution under the assumption that $u\in L^{\infty}((0,T_{\rm max});L^{\frac{n}{2}+}(\Omega))$
Lemma 4.3.
Let $\chi>0$, and suppose that $u\in L^{\infty}\left((0,T_{\rm max});L^{p}(\Omega)\right)$ for some $p>\frac{n}{2}$. Then $u\in L^{\infty}\left((0,T_{\rm max});L^{\infty}(\Omega)\right)$.
Proof.
Thanks to a uniform lower bound for $v$ as established Lemma 2.3, we follow the standard heat semi-group technique (see Lemma 3.4 in [21]) with the assistance of Lemma 2.4(ii) to prove that $u\in L^{\infty}\left((0,T_{\rm max});L^{\infty}(\Omega)\right)$. . ∎
We are now in position to prove our main result.
Proof of Theorem 1.1.
Applying Corollary 4.2, Lemma 4.3, Lemma 2.4 and the extensibility property of solution (2.1) as established in Lemma 2.1 immediately proves the desired result. ∎